A company is considering two short-term projects:

Month A B

0 -$1,200 -$1800

1 $0 $ 400

2 $200 $ 400

3 $300 $ 300

4 $400 $ 500

5 $500 $ 200

6 $600 $ 100

a. Based on the undiscounted payback period, which project is preferred?

b. Calculate the benefit-cost ratio for each alternative (use excel)

c. Based on the benefit-cost ratio analysis method, which project is preferred?

(Show and explain your solution)
Answer #1

a.

Month

A

Cumulative Cash flow

B

Cumulative Cash flow

0

-1200

-1200

-1800

-1800

1

0

-1200

400

-1400

2

200

-1000

400

-1000

3

300

-700

300

-700

4

400

-300

500

-200

5

500

200

200

0

6

600

800

100

100

Project A = 4+(300/400)

= 4.75 months

Project B = 5 months

So, as the un-discounted payback period is less for project A, project A is preferred

b.

Month

A

B

0

-1200

-1800

1

0

400

2

200

400

3

300

300

4

400

500

5

500

200

6

600

100

Benefit

2000

1900

Cost

1200

1800

B:C ratio

=2000/1200=1.67

=1900/1800=1.06

c. Based on B:C ratio, project A is preferred as the value is higher when compared to project B